# Using Natural Language Explanations to Rescale Human Judgments

### v4

@@ 1 Introduction
With the rise in model performance due to large language models (Brown et al., 2020; Ouyang et al., 2022, LLMs), there is a shift towards working on and evaluating more subjective tasks. On tasks like news summarization, LLMs can no longer reliably be judged by automatic metrics Goyal et al. (2022a); Xu et al. (2023); Wang et al. (2023b) and achieve near-human performance Zhang et al. (2023); Zhan et al. (2023). This high performance makes it harder than ever to annotate model outputs for errors Saunders et al. (2022); Dou et al. (2022); Chang et al. (2023); Goyal et al. (2022b); Liu et al. (2023); Chen et al. (2023a), an important ingredient as models are deployed in high-stake scenarios.
@@
This shift towards more subjective tasks and difficulty in identifying errors has also led to a change in the role of human judgments in NLP. Earlier crowdsourcing work on simple labeling tasks explored capturing annotator bias with item-response models Dawid & Skene (1979); Smyth et al. (1994), especially to learn annotator quality Hovy et al. (2013). As the ambit of NLP research has expanded to include more sophisticated downstream tasks, taking into account subjectivity (and thus, inherent disagreement) in annotated data has surfaced as a key direction Plank (2022); Uma et al. (2021); Basile et al. (2021); Nguyen et al. (2016).
While existing work developed label aggregation methods and ways to incorporate different labels Plank (2022), information from the labels alone is still limited. Instead, intricacies in human judgments can be captured by natural language explanations provided during data annotation, which capture more nuanced subjectivity. They enable us to go beyond direct labeling, where annotators can choose different coarse labels for the same reason or the same coarse label for different reasons. The question is, how do we transfer at least some of that expressiveness into numeric expressions that are more friendly for model training and evaluation?
This paper proposes explanation-based rescaling (EBR) that enables us to transfer the expressiveness of natural language explanations into numeric forms. Our key idea is to make the ordinal label space (e.g., a coarse Likert scale) fine-grained, namely a 0-100 scale Kocmi & Federmann (2023), which then enables us to place the initial annotations in it by leveraging natural language explanations in a principled manner. Our approach begins by gathering explanations for each judgment during the annotation process. Next, we leverage an LLM to convert both a Likert judgment and its associated explanation into a numerical rating between 0-100. Crucially, this rescaling is guided by a predefined aspect-based scoring rubric, which can be defined entirely post-annotation, and provides task-specific guidance to the LLM on the placement of labels on the 0-100 scale. We find that an LLM with a rubric produces scores consistent with those that a human produces when using the same rubric.
@@
We consider the task of evaluating LLM-produced answers for document-grounded, non-factoid question answering. We use annotated high-level questions (targeting discourse understanding) from the Inquisitive dataset Ko et al. (2020), supplementing these with freshly collected questions from geographically diverse and recent text sources. We collect outputs from several LLMs, then annotate answers chiefly for answer completeness, with workers giving a Likert judgment and explanation. Through annotator recruitment and filtering, we selected a set of crowdworkers who gave high-quality, reasoned decisions, yet still had differences in opinion. Figure 2 gives an example from our dataset where annotators chose different Likert judgments for information completeness, but the explanations show that some of them were looking at the same factor. Rescaling with a scoring rubric is able to impart fine-grained distinctions based on these explanations, and in particular, works better than rescaling without a rubric.
@@ 2 Background and Task Desiderata
We need datasets that contain human judgments accompanied by informative, well-formed natural language explanations, where the underlying task satisfies the two properties bolded above (strong model performance and the right amount of subjectivity). We found three datasets that are relevant but they are not compatible with these characteristics. SQuALITY (Wang et al., 2022a) includes evaluation of both human and model summaries by other humans. However, this task is very demanding due to the long document length, and we found that the feedback provided in the existing dataset is short and underspecified. Saunders et al. (2022) also have a dataset with human critiques of topic summaries. However, the evaluation criteria are defined loosely, leading to critiques that are too generic. There is no multiply-annotated data for us to reliably perform analysis on the effect of our rescaling methods. Finally, Filighera et al. (2022) releases a dataset for short answer grading with explanations; however their data does not contain annotator information. Note that other existing datasets for explanations in NLP, such as those mentioned in Wiegreffe & Marasovic (2021), do not align with our task requirements, since classification tasks do not involve fine-grained critiquing of outputs.
@@ 3 Collecting QA pairs
We study human judgments and critiques for non-factoid, document-grounded question answering, focusing on judgments of information completeness, i.e., whether machine-generated answers contain all key pieces of information from a document. We use two sources of data for this task. First, we use the questions collected in the inquisitive dataset Ko et al. (2020). Second, following the annotation guidelines from inquisitive, a linguistics student wrote questions on articles focused on recent news events and non-western geographic entities (Hong Kong, Singapore, and India). This reduces leakage about notable prior news events for the LLMs in question. A summary of the two splits is given in Table 5. All text is in English. We call our dataset Inquisitive-Broad. [^1] The label distributions in the two data splits is similar, suggesting that leakage has no noticeable impact. Thus all results are reported on the entire Inquisitive-Broad. See Appendix A for examples of articles for the task.
The questions target high-level comprehension of text, making the answers complex with information distributed across multiple sentences. They contain a mixture of causal questions (e.g., why are they now liberalizing bank laws?), procedural/elaboration questions (e.g., how will the exceptions to the verification concept be taken care of?), background information questions (e.g., what is the main focus of this movement?), and instantiation questions (e.g., which groups were the human rights activists working on behalf of?Ko et al. (2020), all of which require discourse-level processes and reasoning.
@@
This task meets our desiderata by striking a balance between subjectivity and objectivity. There are divergent opinions about how much information should be included, but because the correct answer is grounded in the article, it should be relatively easy to judge if key information is missing. This contrasts with other long-form question answering tasks like ELI5 Fan et al. (2019), where the lack of grounding makes it difficult to judge information completeness and requires subject matter experts for each question Xu et al. (2023).
@@ 4 Human Evaluation of QA pairs
Using Mechanical Turk, we enlist 8 qualified crowdworkers to collect human evaluation of the question-answer pairs in our dataset. Each instance is evaluated by 5 crowdworkers. The study leads to a dataset of 12.6k annotations consisting of discrete labels along with natural language explanations.
@@ 4.2 Human Label Variation
Figure 3: The proposed rescaling prompt for GPT-4. The rescaling prompt takes as input the aspect definition (eg: completeness is defined in the prompt above), a scoring scale (0-100), the task input (the article), human judgment (Likert rating, natural language explanation and missing sentences) and a scoring rubric.
@@ 5 Explanation-Based Rescaling (EBR)
In this section we propose a rescaling method that captures natural language explanations by incorporating them with the Likert labels. We take inspiration from the scoring mechanism presented in Kocmi & Federmann (2023) for machine translation evaluation, which validated the effectiveness of a 0-100 scale compared to other options. However, instead of using an LLM to directly evaluate the task, we use a more detailed scoring rubric.
@@
Assume we have a collection of items $x_{1},\ldots,x_{n}$ which are being rated by annotators $a_{1},\ldots,a_{m}$ in a sparse manner i.e., not every annotator rates every item. Each annotator assigns an item a discrete rating $r_{ki}$ and writes an explanation $e_{ki}$, where $k$ refers to the item and $i$ refers to the annotator. We rescale annotator judgment as following: $s^{\prime}_{ki}=f(r_{ki},e_{ki})$.
Details about the rubric creation process are given in Appendix C. We also explore rescaling variations without rubric. Variants of the rescaling prompt are given in Appendix D.1 and we report our results in Section 7.1. Note, we do not rescale instances that get extreme labels i.e., complete and missing all, since particularly complete explanations did not show complex explanations in most cases.
@@
    Scoring scale: defines the numeric scale to which the human judgment needs to be mapped (0-100 for our task)
@@
    Task input: helps contextualize the natural language feedback (the article, in our case)
@@ 6.2 Metrics
We determine how well our rescaling method does by comparing the rescaled values with references scores and also by comparing inter-annotator agreement before and after rescaling. We use Mean Absolute Error (MAE) and Kendall’s $\tau$ correlation Kendall (1948).
@@
We also compute a ranking (with ties) of the data instances $M(a_{i})$ according to annotator $a_{i}$. We compute pairwise correlations between two annotators’ rankings $\tau(M_{a_{i}},M_{a_{j}})$. We then compute an aggregate correlation across all pairs of annotators: $\frac{1}{{|A|\choose 2}}\sum_{a_{i},a_{j}\in A\times A,a_{i}\neq a_{j}}\tau(M_{a_{i}},M_{a_{j}})$, where $A$ is the set of annotators.
@@ 6.3 Baselines
Static rescaling maps the Likert ratings to a numeric scale by spacing them out at equal intervals, such that, complete maps to 100, missing minor is a 70, missing major is a 30 and missing all is 0. This mapping does not use any fine-grained information from the explanations.
Avg EBR rescaling leverages the explanations to come up with a mapping for the four Likert labels to a 0-100 scale. To find this mapping, we first use EBR to rescale human judgment in the dataset. We then average all numeric scores under each of the labels. This method allows us to incorporate explanations while keeping the original level of granularity. This also maps labels at a calibrated interval from each other instead of having them at equal intervals (as in the Static rescaling method). We get the following mapping for each label: complete – 100, missing minor – 78.6, missing major – 50.9, missing all – 0.0.
@@
7 Results
7.1 Rescaling Results
Are humans able to do the proposed rescaling consistently?
Table 3 shows how our proposed rescaling using GPT-4 compares to reference scores (on 145 instances mentioned in Section 6.1). We look at Kendall’s $\tau$ and MAE between reference scores (average of expert rescaled scores) and EBR scores. We also compare EBR with the baseline methods defined in Section 6.3. Again, we specifically look at missing minor and missing major label categories, since they are the main source of subjectivity and have more nuanced explanations.
@@ 7.2 Annotator Alignment
Table 4 shows that our method is able to rescale human judgment to a much more fine-grained scale without impacting agreement i.e. we are able to go more granular in our scoring without impacting the inter-annotator correlation.
@@
How does rescaling impact subjectivity?
We emphasize that our goal is not to “smooth out” subjectivity in the annotation. Rather, taking natural language explanations into account produces more calibrated and nuanced view that preserves inherent subjectivity while aligning differences between annotators when they actually agree.
@@ 8 Related Work
Kocmi & Federmann (2023) present a scoring mechanism for machine translation that we take inspiration from in this work. Other work has investigated giving prompts directly to LLMs instead of to annotators Chiang & Lee (2023); Wang et al. (2023a); Chen et al. (2023b). Similarly, He et al. (2023); Gilardi et al. (2023); Törnberg (2023); Zhu et al. (2023) showed that GPT-4 can outperform average crowd workers in several annotation tasks including BoolQ (Clark et al., 2019), Twitter content moderation (Alizadeh et al., 2022), sentiment analysis (Rosenthal et al., 2017), and more.  Wang et al. (2022b) showed that GPT-3.5 can generate near expert-human instructions to align LLMs’ behavior to instructions. Previous work has also explored the use of a defined criteria for model-based fine-grained error localization and rationale generation  Kim et al. (2023); Jiang et al. (2023). Our work is assumes that some tasks on the frontiers of LLM capabilities will always need human judgments; our procedure aims to augment and improve the capabilities of the group of crowdworkers.
@@
Another method for rescaling annotations is the calibrated sigma method Weijters et al. (2016). However, this method does not change the ranking of examples, only rescales their scores for each annotator while preserving the ranking. We therefore do not compare to it here. Another line of work Hovy et al. (2013); Gordon et al. (2021) focuses on modeling annotators to build better consensus judgments, but does not address aligning pairs of annotators. Finally, Ethayarajh & Jurafsky (2022) present a new protocol for NLG evaluation in which annotators give judgments in the form of probabilities over sets. This work also demonstrates the flaws in Likert judgments and pursues an approach to improving them orthogonal to ours.
@@
Natural language explanations (NLE) have been shown to be effective in improving model performance when used as additional features (Wu & Mooney, 2019; Murty et al., 2020; Liang et al., 2020), explaining model decisions (Camburu et al., 2018; Narang et al., 2020; Hase et al., 2020), boosting the performance of in-context learning of LLMs on a variety of reasoning tasks (Nye et al., 2021; Wei et al., 2022; Lampinen et al., 2022). Prior work also demonstrated that natural language explanations help us holistically understand annotator perspectives in complex tasks Ferracane et al. (2021); Goyal et al. (2022a), improve annotation quality Alonso (2013) and aid adjudication Filighera et al. (2022). Our work uses a similar idea but focuses on rescaling human annotation specifically, which is not a task addressed in this prior work.

### v6

@@ 1 Introduction
With the rise in model performance due to large language models (Brown et al., 2020; Ouyang et al., 2022, LLMs), there is a shift towards working on and evaluating more subjective tasks. On tasks like news summarization, LLMs can no longer reliably be judged by automatic metrics (Goyal et al., 2022a; Xu et al., 2023; Wang et al., 2023b) and achieve near-human performance (Zhang et al., 2023; Zhan et al., 2023). This high performance makes it harder than ever to annotate model outputs for errors (Saunders et al., 2022; Dou et al., 2022; Chang et al., 2023; Goyal et al., 2022b; Liu et al., 2023; Chen et al., 2024), an important ingredient as models are deployed in high-stake scenarios.
@@
This shift towards more subjective tasks and difficulty in identifying errors has also led to a change in the role of human judgments in NLP. Earlier crowdsourcing work on simple labeling tasks explored capturing annotator bias with item-response models (Dawid & Skene, 1979; Smyth et al., 1994), especially to learn annotator quality (Hovy et al., 2013). As the ambit of NLP research has expanded to include more sophisticated downstream tasks, taking into account subjectivity (and thus, inherent disagreement) in annotated data has surfaced as a key direction (Plank, 2022; Uma et al., 2021; Basile et al., 2021; Nguyen et al., 2016).
While existing work developed label aggregation methods and ways to incorporate different labels (Plank, 2022), information from the labels alone is still limited. Instead, intricacies in human judgments can be captured by natural language explanations provided during data annotation, which capture more nuanced subjectivity. They enable us to go beyond direct labeling, where annotators can choose different coarse labels for the same reason or the same coarse label for different reasons. The question is, how do we transfer at least some of that expressiveness into numeric expressions that are more friendly for model training and evaluation?
This paper proposes explanation-based rescaling (EBR) that enables us to transfer the expressiveness of natural language explanations into numeric forms. Our key idea is to make the ordinal label space (e.g., a coarse Likert scale) fine-grained, namely a 0-100 scale (Kocmi & Federmann, 2023), which then enables us to place the initial annotations in it by leveraging natural language explanations in a principled manner. Our approach begins by gathering explanations for each judgment during the annotation process. Next, we leverage an LLM to convert both a Likert judgment and its associated explanation into a numerical rating between 0-100. Crucially, this rescaling is guided by a predefined aspect-based scoring rubric, which can be defined entirely post-annotation, and provides task-specific guidance to the LLM on the placement of labels on the 0-100 scale. We find that an LLM with a rubric produces scores consistent with those that a human produces when using the same rubric.
@@
We consider the task of evaluating LLM-produced answers for document-grounded, non-factoid question answering. We use annotated high-level questions (targeting discourse understanding) from the Inquisitive dataset (Ko et al., 2020), supplementing these with freshly collected questions from geographically diverse and recent text sources. We collect outputs from several LLMs, then annotate answers chiefly for answer completeness, with workers giving a Likert judgment and explanation. Through annotator recruitment and filtering, we selected a set of crowdworkers who gave high-quality, reasoned decisions, yet still had differences in opinion. Figure 2 gives an example from our dataset where annotators chose different Likert judgments for information completeness, but the explanations show that some of them were looking at the same factor. Rescaling with a scoring rubric is able to impart fine-grained distinctions based on these explanations, and in particular, works better than rescaling without a rubric.
@@ 2 Background and Task Desiderata
We need datasets that contain human judgments accompanied by informative, well-formed natural language explanations, where the underlying task satisfies the two properties bolded above (strong model performance and the right amount of subjectivity). We found three datasets that are relevant but they are not compatible with these characteristics. SQuALITY (Wang et al., 2022) includes evaluation of both human and model summaries by other humans. However, this task is very demanding due to the long document length, and we found that the feedback provided in the existing dataset is short and underspecified. Saunders et al. (2022) also have a dataset with human critiques of topic summaries. However, the evaluation criteria are defined loosely, leading to critiques that are too generic. There is no multiply-annotated data for us to reliably perform analysis on the effect of our rescaling methods. Finally, Filighera et al. (2022) releases a dataset for short answer grading with explanations; however their data does not contain annotator information. Note that other existing datasets for explanations in NLP, such as those mentioned in Wiegreffe & Marasovic (2021), do not align with our task requirements, since classification tasks do not involve fine-grained critiquing of outputs.
@@ 3 Collecting QA pairs
We study human judgments and critiques for non-factoid, document-grounded question answering, focusing on judgments of information completeness, i.e., whether machine-generated answers contain all key pieces of information from a document. We use two sources of data for this task. First, we use the questions collected in the inquisitive dataset (Ko et al., 2020). Second, following the annotation guidelines from inquisitive, a linguistics student wrote questions on articles focused on recent news events and non-western geographic entities (Hong Kong, Singapore, and India). This reduces leakage about notable prior news events for the LLMs in question. A summary of the two splits is given in Table 5. All text is in English. We call our dataset Inquisitive-Broad. [^1] The label distributions in the two data splits is similar, suggesting that leakage has no noticeable impact. Thus all results are reported on the entire Inquisitive-Broad. See Appendix A for examples of articles for the task.
The questions target high-level comprehension of text, making the answers complex with information distributed across multiple sentences. They contain a mixture of causal questions (e.g., why are they now liberalizing bank laws?), procedural/elaboration questions (e.g., how will the exceptions to the verification concept be taken care of?), background information questions (e.g., what is the main focus of this movement?), and instantiation questions (e.g., which groups were the human rights activists working on behalf of?(Ko et al., 2020), all of which require discourse-level processes and reasoning.
@@
This task meets our desiderata by striking a balance between subjectivity and objectivity. There are divergent opinions about how much information should be included, but because the correct answer is grounded in the article, it should be relatively easy to judge if key information is missing. This contrasts with other long-form question answering tasks like ELI5 (Fan et al., 2019), where the lack of grounding makes it difficult to judge information completeness and requires subject matter experts for each question (Xu et al., 2023).
@@ 4 Human Evaluation of QA pairs
Using Mechanical Turk, we enlist 8 qualified crowdworkers to conduct human evaluation of the question-answer pairs in our dataset. Each instance is evaluated by 5 crowdworkers. The study leads to a dataset of 12.6k annotations consisting of discrete labels along with natural language explanations.
@@ 4.2 Human Label Variation
Figure 3: The rescaling prompt for GPT-4, which considers the aspect definition (eg: completeness, defined in the prompt above), a scoring scale (0-100), the task input (the article), human judgment (Likert rating, natural language explanation and missing sentences) and a scoring rubric.
@@ 5 Explanation-Based Rescaling (EBR)
In this section we propose a rescaling method that captures natural language explanations by using them to re-evaluate the Likert labels. We take inspiration from the scoring mechanism presented in Kocmi & Federmann (2023) for machine translation evaluation, which validated the effectiveness of a 0-100 scale compared to other options. The key element here is that a chosen scale provides higher granularity than the coarse grained scale on which the initial Likert labels are collected. Instead of using an LLM to directly evaluate the task, we use a more detailed scoring rubric.
@@
Assume we have a collection of items $x_{1},\ldots,x_{n}$ which are being rated by annotators $a_{1},\ldots,a_{m}$ in a sparse manner, i.e., not every annotator rates every item. Each annotator assigns an item a discrete rating $r_{ki}$ and writes an explanation $e_{ki}$, where $k$ refers to the item and $i$ refers to the annotator. We rescale annotator judgment as following: $s^{\prime}_{ki}=f(r_{ki},e_{ki})$.
Details about the rubric creation process are given in Appendix C. We also explore rescaling variations without rubric. Variants of the rescaling prompt are given in Appendix D.1 and we report our results in Section 7.1. Note, we do not rescale instances that get the extreme labels complete and missing all, since in these cases we rarely observed nuanced explanations from which rescaling would be possible.
@@
    Scoring scale: defines the numeric scale to which the human judgment needs to be mapped (0-100 for our task; note that the LLM typically outputs multiples of 5)
@@
    Task input: helps contextualize the natural language feedback (the article)
@@ 6.2 Metrics
We determine how well our rescaling method does by comparing the rescaled values with references scores and also by comparing inter-annotator agreement before and after rescaling. We use Mean Absolute Error (MAE) and Kendall’s $\tau$ correlation (Kendall, 1948).
@@
We also compute pairwise correlations between two annotators’ rankings, which we denote as $\tau(M(r_{ki},e_{ki}),M(r_{kj},e_{kj}))$ for annotators $i$ and $j$, where $k$ is a placeholder index. We then compute an aggregate correlation across all pairs of annotators: $\frac{1}{{|A|\choose 2}}\sum_{a_{i},a_{j}\in A\times A,a_{i}\neq a_{j}}\tau(M(r_{ki},e_{ki}),M(r_{kj},e_{kj}))$, where $A$ is the set of annotators.
@@ 6.3 Baselines
Static rescaling maps the Likert ratings to a static numeric scale by assigning complete to 100, missing minor to 70, missing major to 30 and missing all to 0. This mapping does not use any fine-grained information from the explanations.
Avg EBR rescaling leverages the explanations to come up with a mapping for the four Likert labels to a 0-100 scale. To find this mapping, we first use EBR to rescale human judgment in the dataset. We then average all numeric scores under each of the labels. This method allows us to incorporate explanations while keeping the original level of granularity. This also maps labels at a calibrated interval from each other instead of having them at equal intervals (as in the Static rescaling method). We get the following mapping for each label: complete: 100, missing minor: 78.6, missing major: 50.9, missing all: 0.0.
@@ 7 Results
7.1 Rescaling Results
Are humans able to do the proposed rescaling consistently?
Table 3 shows how our proposed rescaling using GPT-4 compares to reference scores on 145 instances mentioned in Section 6.1. We look at Kendall’s $\tau$ and MAE between reference scores (average of expert rescaled scores) and EBR scores. We also compare EBR with the baseline methods defined in Section 6.3. Again, we specifically look at missing minor and missing major label categories, since they are the main source of subjectivity and have more nuanced explanations.
@@ 7.2 Annotator Alignment
Table 4 shows that our method is able to rescale human judgment to a much more fine-grained scale without impacting agreement; that is, our scoring can be more granular without impacting the inter-annotator correlation.
@@
How does rescaling impact subjectivity?
We emphasize that our goal is not to “smooth out” subjectivity in the annotation. Rather, taking natural language explanations into account produces a more calibrated and nuanced view that preserves inherent subjectivity while aligning differences between annotators when they actually agree.
@@ 8 Related Work
Kocmi & Federmann (2023) present a scoring mechanism for machine translation that we take inspiration from in this work. Other work has investigated giving prompts directly to LLMs instead of to annotators (Chiang & Lee, 2023; Wang et al., 2023a; Chen et al., 2023). Similarly, He et al. (2024); Gilardi et al. (2023); Törnberg (2023); Zhu et al. (2023) showed that GPT-4 can outperform average crowd workers in several annotation tasks including BoolQ (Clark et al., 2019), Twitter content moderation (Alizadeh et al., 2022), sentiment analysis (Rosenthal et al., 2017), and more.  Wang et al. (2023c) showed that GPT-3.5 can generate near expert-human instructions to align LLMs’ behavior to instructions. Previous work has also explored the use of a defined criteria for model-based fine-grained error localization and rationale generation  (Kim et al., 2024; Jiang et al., 2024). Our work is assumes that some tasks on the frontiers of LLM capabilities will always need human judgments; our procedure aims to augment and improve the capabilities of the group of crowdworkers.
@@
Another method for rescaling annotations is the calibrated sigma method (Weijters et al., 2016). However, this method does not change the ranking of examples, only rescales their scores for each annotator while preserving the ranking. We therefore do not compare to it here. Another line of work (Hovy et al., 2013; Gordon et al., 2021) focuses on modeling annotators to build better consensus judgments, but does not address aligning pairs of annotators. Sakaguchi & Van Durme (2018) proposed a method EASL, that adapts the original annotator labels to probabilistic ones based on the labels of all annotators. Finally, Ethayarajh & Jurafsky (2022) present a new protocol for NLG evaluation in which annotators give judgments in the form of probabilities over sets. This work also demonstrates the flaws in Likert judgments and pursues an approach to improving them orthogonal to ours. Consensus building has also been extensively studied in the translation community (Bojar et al., 2016). Graham et al. (2013) showed that using a continuous scale of 0-100 improved inter-annotator correlations on translation tasks.
@@
Natural language explanations (NLE) have been shown to be effective in improving model performance when used as additional features (Wu & Mooney, 2019; Murty et al., 2020; Liang et al., 2020), explaining model decisions (Camburu et al., 2018; Narang et al., 2020; Hase et al., 2020), boosting the performance of in-context learning of LLMs on a variety of reasoning tasks (Nye et al., 2021; Wei et al., 2022; Lampinen et al., 2022). Prior work also demonstrated that natural language explanations help us holistically understand annotator perspectives in complex tasks (Ferracane et al., 2021; Goyal et al., 2022a), improve annotation quality (Alonso, 2013) and aid adjudication (Filighera et al., 2022). Our work uses a similar idea but focuses on rescaling human annotation specifically, which is not a task addressed in this prior work.